OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. m = , P(3, 2) and Q(8,_) -1 1 5

OpenStudy (anonymous):

@Skielerlucas04 I need your help again if u don't mind

OpenStudy (anonymous):

is that the whole problem?

OpenStudy (anonymous):

wait theres a fraction

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. m = -7/2 , T(13, 2), W(_, -5) -11 11 15

OpenStudy (anonymous):

idk this sorry

OpenStudy (anonymous):

that's okay neither do i

OpenStudy (anonymous):

For the 1st question, what is m=?

OpenStudy (anonymous):

Please check the question because one information is missing.

OpenStudy (anonymous):

it was a fraction 3/5

OpenStudy (anonymous):

\[\large m=\frac{ y _{2}-y_{1} }{ x_{2}-x_{1} }\] \[solve~for~y_{2}.\]\[\large \frac{ 3 }{ 5 }=\frac{ y _{2}-2 }{8-3}\]